# SARDP_supp

### v1

@@ 3.4. Theoretical Properties of the CLS
With $\hat{\gamma}_{\text{CLS}}$ plugged in, the consistent estimators for $\Sigma_{k}^{\text{LS}}$ could be obtained as $\widehat{\Sigma}_{k}^{\text{LS}}=\Sigma_{k}^{\text{LS}}(\mathbb{X}^{*},\hat{\gamma}_{\text{CLS}})-\Delta_{k}^{\text{LS}}(\hat{\gamma}_{\text{CLS}})$ $(k=1,2)$. In this way, we could obtain the estimated standard error of $\hat{\gamma}_{\text{CLS}}$ by the estimated asymptotic covariance matrix $(\widehat{\Sigma}_{2}^{\text{LS}})^{-1}\widehat{\Sigma}_{1}^{\text{LS}}(\widehat{\Sigma}_{2}^{\text{LS}})^{-1}$. Using a technique similar to that used in the proof of Lemma 8 in Appendix B, the consistency of $\widehat{\Sigma}_{k}^{\text{LS}}$ can be established.
@@ 4.1. Simulation Models
To demonstrate the finite-sample performance of the proposed methods, we present three simulation examples based on the following generating mechanisms: the network structure $A$, data generation, noise distribution, and privacy-protection level. The proposed estimators (i.e., CLE and CLS) are compared. The network structures are given as follows.
@@
Data Generation. For each node, we generate the true exogenous covariates $\mathbf{x}_{i}=(x_{i1},x_{i2})^{\top}\in\mathbb{R}^{2}$ from a multivariate normal distribution with mean $\mathbf{0}_{2}$ and $\Sigma_{x}=I_{2}$. The corresponding network autoregression coefficient is fixed to be $\rho_{0}=0.2$ and coefficient $\beta_{0}=(0.3,0.3)^{\top}$. The response $\mathbb{Y}^{*}$ is generated based on $\mathbb{Y}^{*}=(I_{N}-\rho_{0}W)^{-1}(\mathbb{X}\beta_{0}+{\mathbb{E}})+{\mathcal{E}}$. Define $X_{j}=(x_{1j},\cdots,x_{nj})^{\top}$ for $1\leq j\leq 2$. To simulate privacy-protected covariates, we fix $\mathbb{X}_{1}=X_{1}\in\mathbb{R}^{N}$ and assume that $\mathbb{X}_{2}=X_{2}\in\mathbb{R}^{N}$ cannot be faithfully observed. Then, $\mathbb{X}_{2}^{*}$ can be generated by $\mathbb{X}_{2}^{*}=\mathbb{X}_{2}+{\mathcal{E}}$ with ${\mathcal{E}}=(\epsilon_{x,i})$, which follows the setting of the noise distribution.
Noise Distribution. We consider that $\epsilon_{x,i}$ follows a normal distribution with mean 00 and $\lambda_{x}^{2}=0.5$. We consider two different distributions to generate $\epsilon_{i}$ independently with mean 00 and $\lambda^{2}=0.5$: (1) a normal distribution, and (2) a $t$-distribution with degree 6. It is remarkable that for the $t$-distribution, each element is divided by $\sqrt{3}$ to make $\mbox{var}(\epsilon_{i})=\lambda^{2}=0.5$. For the noise $e_{i}$, we consider the same cases for $N(0,1)$ and $t$(6). Here, we set the sample size as $N=$500, 1,000, and 2,000 and consider all the network models.
Privacy-Protection Level. To better illustrate the effect of the privacy-protection level, we consider a fixed sample size $N=1,000$ and the dyad independent network-generation model as an example. Different $\epsilon_{i}$s and $\epsilon_{x,i}$s are generated from normal distributions with variances of $\lambda^{2}$ and $\lambda^{2}_{x}$, respectively. We consider the impact of the variations in $\lambda_{x}^{2}$ and $\lambda^{2}$ on the estimation results. Specifically, for fixed $\lambda^{2}=0.5$, we consider $\lambda^{2}_{x}=(0.2,0.5,0.8)$. For fixed $\lambda^{2}_{x}=0.5$, we consider $\lambda^{2}=(0.2,0.5,0.8).$
@@ 4.2. Performance Measurements and Simulation Results
For a better comparison, we focus here on the estimation results of $\rho_{0}$ and $\beta_{0}$. To gauge the finite-sample performance, we use the following metrics. Define $\hat{D}^{(r)}=\{\hat{d}_{j}^{(r)}\}_{j=1}^{q}\in\mathbb{R}^{q}$ as the estimator from the $r$th replication with $q=p+1$. For any $1\leq j\leq q$, the bias can be evaluated as ${\rm Bias}_{j}=|\bar{d_{j}}-d_{j}|$, where $\bar{d_{j}}=R^{-1}\sum_{r}\hat{d}_{j}^{(r)}$, and $d_{j}$ is the $j$th element of the true parameter. The standard error can be estimated using $\widehat{\mbox{SE}}_{j}=R^{-1}\sum_{r}\widehat{\mbox{SE}}_{j}^{(r)}$. Notably, $\widehat{\mbox{SE}}_{j}^{(r)}$ represents the $j$th diagonal element of the estimated asymptotic covariance matrix, which can be computed using (3.3) or (3.4) with $\hat{D}^{(r)}$ plugged in, followed by a correction step, as discussed in Section 3.3 for the CLE and Section 3.4 for the CLS. Define ${\rm SE}_{j}=\{R^{-1}\sum_{r}(\hat{d}_{j}^{(r)}-\bar{d_{j}})^{2}\}^{1/2}$ as the Monte Carlo standard deviation of $\hat{d}_{j}^{(r)}$, and the estimation efficiency of $\widehat{\rm SE}_{j}$ can be evaluated by comparing $\widehat{\mbox{SE}}_{j}$ and ${\rm SE}_{j}$. In addition, for each $\hat{d}_{j}^{(r)}$, a 95% confidence interval could be constructed as $\mbox{CI}_{j}^{(r)}=[\hat{d}_{j}^{(r)}-z_{0.975}N^{-1}\widehat{\mbox{SE}}_{j}^{(r)},\hat{d}_{j}^{(r)}+z_{0.975}N^{-1}\widehat{\mbox{SE}}_{j}^{(r)}]$, where $z_{\alpha}$ is the lower $\alpha$th quantile of the standard normal distribution. The empirical coverage probability is then evaluated as $\mbox{CP}_{j}=R^{-1}\sum_{r=1}^{R}I(\hat{d}_{j}^{(r)}\in\mbox{CI}_{j}^{(r)})$, where $I(\cdot)$ is the indicator function.
@@
Each experiment is replicated 500 times ($R=500$). All simulations are conducted on a Linux server with a 3.60 GHz Intel Core i7-9700K CPU and 16 GB RAM. Due to the similar estimation performances across different noise distributions, we only present the results for the normal distributions of both ${\mathbb{E}}$ and ${\mathcal{E}}$ in Tables 2, and the other results are provided in Appendix D. Moreover, the estimation outcomes for different values of $\lambda^{2}$ and $\lambda_{x}^{2}$ are illustrated in Table 3. We present the Bias, SE, and CP in Tables 2-3. Then, we show the averaged CPU time for deriving CLE and CLS in Figure 1 to evaluate the computation efficiency.
@@
Estimation Performance. From Table 2, we draw the following conclusions. First, the estimation bias is sufficiently small for all sample sizes. Second, as the sample size $N$ increases, Bias, SE, and $\widehat{\mbox{SE}}$ all decrease, which shows the consistency of all three methods. Third, all the CP values are approximately 95% for $\alpha=0.05$ with $\widehat{\mbox{SE}}$ and SE clearly being close to each other. This corroborates the theoretical conclusions in Theorem 2 for the CLE and in Theorem 5 for the CLS. Finally, CLE has a smaller SE than CLS.
Effect of Privacy-Protection Level. Based on the results in Table 3, we can conclude that an increase in the privacy-protection level leads to a larger $\widehat{\rm SE}$ for the estimators. Specifically, when $\lambda_{x}^{2}$ is fixed, increasing $\lambda^{2}$ leads to an increase in $\widehat{\rm SE}$ for all estimators. However, when $\lambda^{2}$ is fixed, as $\lambda_{x}^{2}$ increases, only $\widehat{\rm SE}$ for the estimator corresponding to the unobserved $\mathbb{X}_{2}$ increases, whereas those for the other estimators remain relatively stable.
@@
Computation Time. We conduct the same experiment as previously, but with a normal noise distribution and fixed $\sigma_{0}^{2}=1,$ $\lambda_{x}^{2}=1,$ and $\lambda^{2}=0.5$. Moreover, we change $N$ from 500 to 5,000 for $R=100$ (100 replicates) to further compare the computational efficiency of CLE and CLS. Figure 1 shows the results from a computational perspective. Evidently, as the sample size $N$ increases, the computational time of CLE increases dramatically, whereas that of CLS increases much slower. In practice, a method should be chosen based on the balance between computational time and statistical accuracy.

### v2

@@ 3.4. Theoretical Properties of the CLS
Through Theorem 5, we can conclude that the CLS estimator $\hat{\gamma}_{\text{CLS}}$ is $\sqrt{N}$-consistent, which has the same convergence rate as the CLE. For statistical inference, similar to the discussion for $\hat{\theta}_{\text{CLE}}$, we can also provide consistent estimators for each element in the asymptotic covariance matrix (3.2) based on the observed $\mathbb{X}^{*}$. Rewrite $\Sigma_{k}^{\text{LS}}=\Sigma_{k}^{\text{LS}}(\mathbb{X},\gamma)$ $(k=1,2)$ as functions of $\mathbb{X}$ and $\gamma$, which are expressed in (3.2). Then, we could plug $\mathbb{X}^{*},\hat{\gamma}_{\text{CLS}}$ in to obtain $\Sigma_{k}^{\text{LS}}(\mathbb{X}^{*},\hat{\gamma}_{\text{CLS}})$ for $k=1,2$. As a result, $\Delta_{k}^{\text{LS}}(\gamma)=E\{\Sigma_{k}^{\text{LS}}(\mathbb{X}^{*},\theta)\}-\Sigma_{k}^{\text{LS}}(\mathbb{X},\theta)$ could be calculated. See Appendix A.2 for a detailed expression of $\Delta_{k}^{\text{LS}}(\gamma)$. With $\hat{\gamma}_{\text{CLS}}$ plugged in, the consistent estimators for $\Sigma_{k}^{\text{LS}}$ could be obtained as $\widehat{\Sigma}_{k}^{\text{LS}}=\Sigma_{k}^{\text{LS}}(\mathbb{X}^{*},\hat{\gamma}_{\text{CLS}})-\Delta_{k}^{\text{LS}}(\hat{\gamma}_{\text{CLS}})$ $(k=1,2)$. In this way, we could obtain the estimated standard error of $\hat{\gamma}_{\text{CLS}}$ by the estimated asymptotic covariance matrix $(\widehat{\Sigma}_{2}^{\text{LS}})^{-1}\widehat{\Sigma}_{1}^{\text{LS}}(\widehat{\Sigma}_{2}^{\text{LS}})^{-1}$. Using a technique similar to that used in the proof of Lemma 8 in Appendix B, the consistency of $\widehat{\Sigma}_{k}^{\text{LS}}$ can be established.
3.5. Model Extensions and Related Theoretical Discussions
The previously discussed model (2.2) represents the simplest form of the SAR model. In this subsection, we study several extensions. Specifically, (1) we consider multiple dependent variables $Y$ instead of a single one, extending the SAR model to the multivariate spatial autoregressive model (Zhu et al., 2020, MSAR); (2) we consider networks with different types of nodes, extending the network to a multi-mode network regression model that accommodates varying node types (Huang et al., 2020); (3) we consider perturbations in the network structure as a preliminary attempt, exploring how they impact the robustness and reliability of the privacy-preserving techniques (Lewbel et al., 2024).
Privacy-Protected Multivariate SAR. First, we consider extending the PSAR model to the privacy-protected MSAR model, which can be adapted to model the mutual influence of different dependent variables through the network structure. Suppose there are multivariate responses $\mathbb{Y}=(Y_{ij})\in\mathbb{R}^{N\times q}$ and error matrix $\tilde{{\mathbb{E}}}=(e_{ij})\in\mathbb{R}^{N\times q}$. Denote the network parameters as $D=(d_{j^{\prime}j})\in\mathbb{R}^{q\times q}$ and the covariates parameters as $B=(b_{kj})\in\mathbb{R}^{p\times q}$. The multivariate spatial autoregressive model is defined as $\mathbb{Y}=W\mathbb{Y}D+\mathbb{X}B+\tilde{\mathbb{E}}$. Here, $d_{j_{1},j_{2}}$ ($1\leq j_{1}\neq j_{2}\leq q$) represents the extra-activity effect, measuring network effect between different responses, and $d_{jj}$ ($1\leq j\leq q$) represents the inter-activity effect, measuring network effect within the same response. Further, we define $\mathcal{Y}={\rm vec}(\mathbb{Y})\in\mathbb{R}^{Nq},\mathcal{X}=I_{q}\otimes\mathbb{X}\in\mathbb{R}^{Nq\times pq}$. Then the vector norm of the privacy-protected MSAR is,
$\mathcal{Y}=(D^{\top}\otimes W)\mathcal{Y}+\mathcal{X}{\rm vec}(B)+{\rm vec}(\tilde{\mathbb{E}}),$
where the observed data is given by $\mathcal{Y}^{*}=\mathcal{Y}+\mathcal{E}$, $\mathcal{X}^{*}=(\mathcal{X}_{1},\mathcal{X}_{2}^{*})=(\mathcal{X}_{1},\mathcal{X}_{2}+\mathcal{E}_{x})$, $\mathcal{E}$ and $\mathcal{E}_{x}$ are artificially added independent noises with zero mean and variance $\lambda^{2}$ and $\lambda_{x}^{2}$ correspondingly. In this way, define $\gamma_{M}\in\mathbb{R}^{q^{2}+pq}$ as the parameter vector that incorporates all the information from $D$ and $B$. We construct the objective function and obtain the CLS estimator $\hat{\gamma}_{M,\rm CLS}$ for $\gamma_{M}$ by correcting for the first- and second-order derivatives to deal with observational errors in $\mathcal{X}^{*}$ and $\mathcal{Y}^{*}$. Consequently, we can establish the following corollary in a manner similar to Theorem 5. Detailed notations are provided in Appendix D, and the proof is omitted.
Corollary 1.
(Asymptotic Normality for $\hat{\gamma}_{M,\rm CLS}$ ) Assume that (C1)–(C3) and (C4) hold. We then have $\sqrt{N}(\hat{\gamma}_{M,\rm CLS}-\gamma_{M})\to_{d}N(\mathbf{0}_{q^{2}+pq},(\Sigma_{2}^{M,\rm LS})^{-1}\Sigma_{1}^{M,\rm LS}(\Sigma_{2}^{M,\rm LS})^{-1})$ as $N\to\infty$, where $\Sigma_{1}^{M,\rm LS}$ and $\Sigma_{2}^{M,\rm LS}$ are assumed to be positive definite matrices. See Appendix D for the detailed formula.
Privacy-Protected Multi-Mode SAR. Next, we employ the privacy-protected multi-mode SAR model to measure the network influence among different types of nodes. Consider a multi-mode network with $n_{k}$ nodes in the $k$th group ($1\leq k\leq K$), and $N=\sum_{k}n_{k}$. Define $a_{i_{1}i_{2}}=0$ if $i_{1}$ and $i_{2}$ are in the same group. Let $Y_{k}=(Y_{k,1},\cdots,Y_{k,n_{k}})^{\top}\in\mathbb{R}^{n_{k}}$ $(1\leq k\leq K)$ be the response collected for the $k$th group of nodes. And define $X_{k}=(X_{k,1}^{\top},\cdots,X_{k,n_{k}}^{\top})^{\top}\in\mathbb{R}^{n_{k}\times p_{k}}$ as the exogenous covariates for the $k$th group. Furthermore, define $\varepsilon_{k}=(\varepsilon_{k,1},\cdots,\varepsilon_{k,n_{k}})^{\top}\in\mathbb{R}^{n_{k}}$ as the noise vector. The multi-mode autoregressive model can be defined as
$Y_{k_{1}}=\sum_{k_{2}\neq k_{1}}\rho_{k_{1}k_{2}}W_{k_{1}k_{2}}Y_{k_{2}}+X_{k_{1}}\beta_{k_{1}}+\varepsilon_{k_{1}},$
where $\rho_{k_{1}k_{2}}$ ($1\leq k_{1},k_{2}\leq K,k_{1}\neq k_{2}$) are the cross-mode effects, $\beta_{k_{1}}$ are the parameters corresponding to $X_{k_{1}}$. Similarly, define $\gamma_{T}\in\mathbb{R}^{K(K-1)+\sum_{k}p_{k}}$ as the parameter vector that encompasses all the information from $\rho_{k_{1}k_{2}}$ and $\beta_{k_{1}}$. We can construct the least-square type objective function and obtain the CLS estimator $\hat{\gamma}_{T,\rm CLS}$ for $\gamma_{T}$ based on corrected derivatives to deal with observational errors in $\mathbb{Y}^{*}$ and $\mathbb{X}^{*}$. Using this approach, the following corollary can be established for a two-mode network as an example. Specifically, we observe $\mathbb{Y}^{*},\mathbb{X}^{*}$ with $\mathbb{Y}^{*}=(Y_{1}^{\top},Y_{2}^{\top})^{\top}+\mathcal{E}$, $\mathbb{X}^{*}=[\mathbb{X}^{*}_{1},0_{n_{1}\times p_{2}};0_{n_{2}\times p_{1}},\mathbb{X}^{*}_{2}]^{\top}$, $\mathbb{X}^{*}_{1}=(\mathbb{X}_{11},\mathbb{X}^{*}_{12})=(\mathbb{X}_{11},\mathbb{X}_{12}+\mathcal{E}_{x_{1}})$, $\mathbb{X}^{*}_{2}=(\mathbb{X}_{21},\mathbb{X}^{*}_{22})=(\mathbb{X}_{21},\mathbb{X}_{22}+\mathcal{E}_{x_{2}})$, and $\mathcal{E}\in\mathbb{R}^{n_{1}+n_{2}}$, $\mathcal{E}_{x_{1}}\in\mathbb{R}^{n_{1}\times p_{12}}$, $\mathcal{E}_{x_{2}}\in\mathbb{R}^{n_{2}\times p_{22}}$ are artificially added noises with variance $\lambda^{2},\lambda_{x}^{2}$ respectively, $p_{12}$ and $p_{22}$ represent the dimensions of the covariates requiring privacy protection for the two types of nodes. Detailed expressions are provided in Appendix D and the proof is omitted.
Corollary 2.
(Asymptotic Normality of CLS in $\hat{\gamma}_{T,\rm CLS}$) Assume that (C1)–(C3) and (C4) hold. We then have $\sqrt{N}(\hat{\gamma}_{T,\rm CLS}-\gamma_{T})\to_{d}N(\mathbf{0}_{K(K-1)+p_{1}+p_{2}},(\Sigma_{2}^{T,\rm LS})^{-1}\Sigma_{1}^{T,\rm LS}(\Sigma_{2}^{T,\rm LS})^{-1})$ as $N\to\infty$, where $\Sigma_{1}^{T,\rm LS}$ and $\Sigma_{2}^{T,\rm LS}$ are assumed to be positive definite matrices. See Appendix D for the detailed formula.
Privacy-Protected SAR with Perturbed Network. As a further discussion of the PSAR model, we consider randomly perturbing the edges in the network to protect the privacy of the network structure. The true network is represented by $A$, and we define the noise-perturbed adjacency matrix as $A^{*}$. In this way, we obtain the observed weighting matrix $W^{*}$, which is normalized from $A^{*}$. Consequently, we calculate the corrected least squares estimator $\hat{\gamma}^{*}_{\rm CLS}$ based on $W^{*}$ instead of $\hat{\gamma}_{\rm CLS}$ based on $W$. However, we cannot allow a large number of edges in the network to change, as excessive noise would prevent us from obtaining consistent parameter estimates. To address this issue, and inspired by Lewbel et al. (2024), we propose the following constraint condition (C5). Based on (C5), we establish the theoretical properties of the CLS under the condition of perturbed network edges. This corollary indicates that once condition (C5) is satisfied, we can still ensure the consistency and asymptotic normality of the CLS estimator.
(C5) (Perturbed Network) Assume $\sum_{i}\sum_{j}E(|A^{*}_{ij}-A_{ij}|)=O(N^{s})$ for some positive constant $s<1/2$, and assume that $W^{*}$ is uniformly bounded in both row and column sums in probability.
Corollary 3.
(Asymptotic Normality of CLS For Perturbed Network) Assume conditions (C1)–(C3), (C4), and (C5) hold. As $N\to\infty$, we then have $\sqrt{N}(\hat{\gamma}^{*}_{\rm CLS}-\gamma_{0})\to_{d}N(\mathbf{0}_{p+1},$$(\Sigma_{2}^{\rm LS})^{-1}\Sigma_{1}^{\rm LS}(\Sigma_{2}^{\rm LS})^{-1}),$ where $\Sigma_{1}^{\rm LS}$ and $\Sigma_{2}^{\rm LS}$ are positive definite matrices defined in Theorem 5. See Appendix D for the detailed formula.
@@ 4.1. Simulation Models
To demonstrate the finite-sample performance of the proposed methods, we present three simulation examples based on the following generating mechanisms: the network structure $A$, data generation, noise distribution, and privacy-protection level. The classical QMLE (treating the observed variables as the true ones) and the proposed estimators (i.e., CLE and CLS) are compared. The network structures are given as follows.
@@
Data Generation. For each node, we generate the true exogenous covariates $\mathbf{x}_{i}=(x_{i1},x_{i2})^{\top}\in\mathbb{R}^{2}$ from a multivariate normal distribution with mean $\mathbf{0}_{2}$ and $\Sigma_{x}=I_{2}$. The corresponding network autoregression coefficient is fixed to be $\rho_{0}=0.2$ and coefficient $\beta_{0}=(0.3,0.3)^{\top}$. The response $\mathbb{Y}^{*}$ is generated based on $\mathbb{Y}^{*}=(I_{N}-\rho_{0}W)^{-1}(\mathbb{X}\beta_{0}+{\mathbb{E}})+{\mathcal{E}}$. Define $X_{j}=(x_{1j},\cdots,x_{nj})^{\top}$ for $1\leq j\leq 2$. To simulate privacy-protected covariates, we fix $\mathbb{X}_{1}=X_{1}\in\mathbb{R}^{N}$ and assume that $\mathbb{X}_{2}=X_{2}\in\mathbb{R}^{N}$ cannot be faithfully observed. Then, $\mathbb{X}_{2}^{*}$ can be generated by $\mathbb{X}_{2}^{*}=\mathbb{X}_{2}+{\mathcal{E}}$ with ${\mathcal{E}}=(\varepsilon_{x,i})$, which follows the setting of the noise distribution.
Noise Distribution. We consider that $\varepsilon_{x,i}$ follows a normal distribution with mean 00 and $\lambda_{x}^{2}=0.5$. We consider two different distributions to generate $\varepsilon_{i}$ independently with mean 00 and $\lambda^{2}=0.5$: (1) a normal distribution, and (2) a $t$-distribution with degree 6. It is notable that for the $t$-distribution, each element is divided by $\sqrt{3}$ to make $\mbox{var}(\varepsilon_{i})=\lambda^{2}=0.5$. For the noise $e_{i}$, we consider the same cases for $N(0,1)$ and $t(6)$. Here, we set the sample size as $N=$500, 1,000, and 2,000 and consider all the network models.
Privacy-Protection Level. To better illustrate the effect of the privacy-protection level, we consider a fixed sample size $N=1,000$ and the dyad independent network-generation model as an example. Different $\varepsilon_{i}$s and $\varepsilon_{x,i}$s are generated from normal distributions with variances of $\lambda^{2}$ and $\lambda^{2}_{x}$, respectively. We consider the impact of the variations in $\lambda_{x}^{2}$ and $\lambda^{2}$ on the estimation results. Specifically, for fixed $\lambda^{2}=0.5$, we consider $\lambda^{2}_{x}=(0.2,0.5,0.8)$. For fixed $\lambda^{2}_{x}=0.5$, we consider $\lambda^{2}=(0.2,0.5,0.8).$
@@ 4.2. Performance Measurements and Simulation Results
For a better comparison, we focus here on the estimation results of $\rho_{0}$ and $\beta_{0}$. To gauge the finite-sample performance, we use the following metrics. Define $\hat{D}^{(r)}=\{\hat{d}_{j}^{(r)}\}_{j=1}^{q}\in\mathbb{R}^{q}$ as the estimator from the $r$th replication with $q=p+1$. For any $1\leq j\leq q$, the bias can be evaluated as ${\rm Bias}_{j}=|\bar{d_{j}}-d_{j}|$, where $\bar{d_{j}}=R^{-1}\sum_{r}\hat{d}_{j}^{(r)}$, and $d_{j}$ is the $j$th element of the true parameter. The standard error can be estimated using $\widehat{\mbox{SE}}_{j}=R^{-1}\sum_{r}\widehat{\mbox{SE}}_{j}^{(r)}$. Notably, $\widehat{\mbox{SE}}_{j}^{(r)}$ represents the $j$th diagonal element of the estimated asymptotic covariance matrix, which can be computed using (3.1) or (3.2) with $\hat{D}^{(r)}$ plugged in, followed by a correction step, as discussed in Section 3.3 for the CLE and Section 3.4 for the CLS. Define ${\rm SE}_{j}=\{R^{-1}\sum_{r}(\hat{d}_{j}^{(r)}-\bar{d_{j}})^{2}\}^{1/2}$ as the Monte Carlo standard deviation of $\hat{d}_{j}^{(r)}$, and the estimation efficiency of $\widehat{\rm SE}_{j}$ can be evaluated by comparing $\widehat{\mbox{SE}}_{j}$ and ${\rm SE}_{j}$. In addition, for each $\hat{d}_{j}^{(r)}$, a 95% confidence interval can be constructed as $\mbox{CI}_{j}^{(r)}=[\hat{d}_{j}^{(r)}-z_{0.975}N^{-1}\widehat{\mbox{SE}}_{j}^{(r)},\hat{d}_{j}^{(r)}+z_{0.975}N^{-1}\widehat{\mbox{SE}}_{j}^{(r)}]$, where $z_{\alpha}$ is the lower $\alpha$th quantile of the standard normal distribution. The empirical coverage probability is then evaluated as $\mbox{CP}_{j}=R^{-1}\sum_{r=1}^{R}I(\hat{d}_{j}^{(r)}\in\mbox{CI}_{j}^{(r)})$, where $I(\cdot)$ is the indicator function.
@@
Each experiment is replicated 500 times ($R=500$). All simulations are conducted on a Linux server with a 3.60 GHz Intel Core i7-9700K CPU and 16 GB RAM. Due to the similar estimation performances across different noise distributions, we only present the results for the normal distributions of both ${\mathbb{E}}$ and ${\mathcal{E}}$ in Tables 2, and the other results are provided in Appendix E. Moreover, the estimation outcomes for different values of $\lambda^{2}$ and $\lambda_{x}^{2}$ are illustrated in Table 3. We present the Bias, SE, and CP in Tables 2-3. Then, we show the averaged CPU time for deriving CLE and CLS in Figure 1 to evaluate the computation efficiency.
@@
Estimation Performance. From Table 2, we can see that the classical QMLE is seriously biased as expected. Thus we focus on the performance of the proposed estimator CLE and CLS. We draw the following conclusions. First, the estimation bias is sufficiently small for all sample sizes and both the methods. Second, as the sample size $N$ increases, Bias, SE, and $\widehat{\mbox{SE}}$ all decrease, which shows the consistency of both methods. Third, all the CP values are approximately 95% for $\alpha=0.05$ with $\widehat{\mbox{SE}}$ and SE clearly being close to each other. This corroborates the theoretical conclusions in Theorem 2 for the CLE and in Theorem 5 for the CLS. Finally, CLE has a smaller SE than CLS.
@@
Effect of Privacy-Protection Level. From Table 3, we can see the bias of $\rho$ or $\beta$ of the classical QMLE increases as the noise variance $\lambda^{2}$ or $\lambda_{x}^{2}$ increases, respectively. Based on the results of CLE and CLS in Table 3, we can conclude that an increase in the privacy-protection level leads to a larger $\widehat{\rm SE}$ for the estimators. Specifically, when $\lambda_{x}^{2}$ is fixed, increasing $\lambda^{2}$ leads to an increase in $\widehat{\rm SE}$ for all estimators. However, when $\lambda^{2}$ is fixed, as $\lambda_{x}^{2}$ increases, only $\widehat{\rm SE}$ for the estimator corresponding to the unobserved $\mathbb{X}_{2}$ increases, whereas those for the other estimators remain relatively stable.
Computation Time. We conduct the same experiment as previously, but with a normal noise distribution and fixed $\sigma_{0}^{2}=1,$ $\lambda_{x}^{2}=1,$ and $\lambda^{2}=0.5$. Moreover, we change $N$ from 500 to 5,000 for $R=100$ (100 replicates) to further compare the computational efficiency of CLE and CLS. Figure 1 shows the results from a computational perspective. Evidently, as the sample size $N$ increases, the computational time of CLE increases dramatically, whereas that of CLS increases much more slowly. In practice, a method should be chosen based on the balance between computational time and statistical accuracy.